# Leveraging Nested MLMC for Sequential Neural Posterior Estimation with Intractable Likelihoods

### v1

@@ Abstract
Sequential neural posterior estimation (SNPE) techniques have been recently proposed for dealing with simulation-based models with intractable likelihoods. They are devoted to learning the posterior from adaptively proposed simulations using neural network-based conditional density estimators. As a SNPE technique, the automatic posterior transformation (APT) method proposed by Greenberg et al. (2019) performs notably and scales to high dimensional data. However, the APT method bears the computation of an expectation of the logarithm of an intractable normalizing constant, i.e., a nested expectation. Although atomic APT was proposed to solve this by discretizing the normalizing constant, it remains challenging to analyze the convergence of learning. In this paper, we propose a nested APT method to estimate the involved nested expectation instead. This facilitates establishing the convergence analysis. Since the nested estimators for the loss function and its gradient are biased, we make use of unbiased multi-level Monte Carlo (MLMC) estimators for debiasing. To further reduce the excessive variance of the unbiased estimators, this paper also develops some truncated MLMC estimators by taking account of the trade-off between the bias and the average cost. Numerical experiments for approximating complex posteriors with multimodal in moderate dimensions are provided.
@@ 1 Introduction
Simulator-based models are widely used across various scientific disciplines, including neuroscience [47], physics [7, 24], biology [10, 29, 39, 46], and inverse graphics [52]. These models serve as crucial tools for describing and comprehending investigated processes based on observed data. However, when applying traditional Bayesian inference to simulator-based models, challenges arise, such as the intractable likelihood function $p(x|\theta)$ and the computational expense associated with running the simulator.
@@
Posterior density estimation approaches approximate the posterior of interest $p(\theta|x_{o})$ with a family of density estimators $q_{\phi}(\theta)$, where $\phi$ is the parameters of the density estimators. Optimization-based approaches are widely used in these methods, and the Kullback-Leibler (KL) divergence between $p(\theta|x)$ and $q_{\phi}(\theta)$, which measures the differences between two densities, is commonly chosen as the loss function. Variational Bayes (VB), as a computationally optimization-based effective method for approximating the posterior distribution of a Bayesian problem, is widely used. In the likelihood-free context, Tran et al., [56] developed a new VB method with an intractable likelihood, while He et al., [30] proposed an unbiased VB method based on nested MLMC. However, these method tend to fail in cases where simulations are expensive, as the nested estimation in these methods requires additional simulation procedures.
Out of this reason, there has been a growing interest in employing neural networks to represent probability density recently, particularly normalizing flows [38]. When the inference problem is focused solely on the observation $x_{o}$, the data efficiency can be improved using the sequential training schemes proposed in the sequential neural posterior estimation (SNPE) methods [25, 42, 48]. In this approach, model parameters are drawn from a proposal distribution that is more informative about $x_{o}$ compared to the prior distribution. However, SNPE requires a modification of the loss function compared to neural posterior estimation (NPE) in order to ensure that the neural networks approximates the true posterior $p(\theta|x)$. Many approaches have been developed to address this issue.
Among the SNPE methods, automatic posterior transformation (APT) requires the computation of the expectation of the logarithm of an intractable normalizing constant, which is a nested expectation. They then propose to use the atomic APT to discretize the normalizing constant. However, it remains challenging to utilize the existing analysis techniques [6] to perform convergence analysis to the best of our knowledge. To address this limitation, we propose the nested APT method, an method enable the convergence analysis without scarifying the performance. We conduct a thorough analysis of the bias and variance of the loss estimators, moreover, we provide a convergence result for this biased estimator in the case of stochastic gradient descent (SGD). To evaluate the effectiveness of the nested APT, we conduct a series of comprehensive numerical experiments.
@@
The remainder of this paper is organized as follows. In Section 2, we propose the nested APT method and provide insights into possible failures when directly using the nested APT. Moreover, we propose a strategy to address this failure. We also present the experimental results and analyze them in this section. In Section 3, we introduce the basics of MLMC methods, including their formulation and theoretical analysis of the random sequence utilized in MLMC methods. Equipped with these tools, we develop both unbiased MLMC methods and truncated MLMC methods and analyze the order of variance for the losses and gradients, as well as the average cost. In Section 4, we provide convergence analysis for nested APT and truncated MLMC methods in the case of SGD. In Section 5, we conduct a series of numerical experiments on benchmark tasks. The results of these experiments are presented and analyzed in this section. Finally, we conclude this paper with some remarks in Section 6.
@@ 2.1 Problem formulation
Let $p(\theta)$ represent the prior distribution of the model parameter of interest. Given an observed sample $x_{o}$, our objective is to perform inference on the posterior distribution $p(\theta|x_{o})\propto p(\theta)p(x_{o}|\theta)$. However, in many cases, the likelihood function $p(x|\theta)$ either lacks an explicit expression or presents challenges in direct evaluation. Instead, it can be expressed using a ‘simulator’ approach, wherein, given a fixed model parameter $\theta$, we can sample $x$ from $p(x|\theta)$.
@@
which is not less than 00 by using Jensen’s inequality and attains its minima when $q(\theta)$ agrees with $p(\theta)$, making it suitable as a loss function. When the likelihood $p(x|\theta)$ is tractable, one can directly approximate $p(\theta|x_{o})$ by minimizing the KL divergence between the target distribution and the proposed estimator $q(\theta)$ within a certain family of distributions. In the likelihood-free context, this can be viewed as a problem of conditional density estimation. Within this framework, a conditional density estimator $q_{F(x,\phi)}(\theta)$ based on a neural network [48, 25], is utilized to approximate $p(\theta|x)$ over the admissible set of tuning parameter $\phi\in\Phi$. To this end, we focus on minimizing the following average KL divergence under the marginal distribution $p(x)=\int p(\theta)p(x|\theta)\mathrm{d}\theta$
@@
$:=\mathcal{L}(\phi)+\iint p(\theta,x)\log p(\theta|x)\mathrm{d}x\mathrm{d}\theta,$
@@
$\mathcal{L}(\phi):=-\mathbb{E}_{p(\theta,x)}\left[\log q_{F(x,\phi)}(\theta)\right],$ (1)
is used as the loss function. However, since $\mathcal{L}(\phi)$ is intractable, we use its empirical estimator instead
^$\hat{\mathcal{L}}_{N}(\phi)=-\frac{1}{N}\sum_{i=1}^{N}\log q_{F(x_{i},\phi)}(\theta_{i}),$ (2)
@@
Since our aim is to conduct conditional density estimation at $x_{o}$, it is essential to utilize a proposal distribution $\tilde{p}(\theta)$ that provides more informative priors regarding $x_{o}$ in comparison to the prior distribution $p(\theta)$. After initializing $\tilde{p}(\theta)$ as $p(\theta)$, we then want the approximation of $p(\theta|x_{o})$ to serve as a good proposal in the following simulation. This conditional density estimation with adaptively chosen proposal is called sequential neural posterior estimation (SNPE). However, after replacing $p(\theta,x)$ with $\tilde{p}(\theta,x)=\tilde{p}(\theta)p(x|\theta)$ in (1), it is observed that $q_{F(x,\phi)}(\theta)$ approximates the proposal posterior:
$\tilde{p}(\theta|x)=p(\theta|x)\frac{\tilde{p}(\theta)p(x)}{p(\theta)\tilde{p}(x)},$ (3)
where $\tilde{p}(x)=\int\tilde{p}(\theta)p(x|\theta)\mathrm{d}\theta$. Hence, we need to adjust loss function $\mathcal{L}(\phi)$ to make $q_{F(x,\phi)}(\theta)$ approximate the true posterior $p(\theta|x)$.
@@
$\mathbb{E}_{\tilde{p}(\theta,x)}\left[-\log\tilde{q}_{F(x,\phi)}(\theta)\right],$ (4)
@@
$\tilde{q}_{F(x,\phi)}(\theta)$ $=q_{F(x,\phi)}(\theta)\frac{\tilde{p}(\theta)}{p(\theta)}\frac{1}{Z(x,\phi)},$ (5)
$Z(x,\phi)$ $=\int\frac{q_{F(x,\phi)}(\theta^{\prime})}{p(\theta^{\prime})}\tilde{p}(\theta^{\prime})\mathrm{d}\theta^{\prime},$
$Z(x,\phi)$ here denotes the normalizing constant. Proposition 1 in [48] shows that if $q_{F(x,\phi)}(\theta)$ is expressive enough that $\tilde{q}_{F(x,\phi^{*})}(\theta)=\tilde{q}(\theta)$ for some parameter $\phi^{*}$, then $q_{F(x,\phi^{*})}(\theta)=q(\theta|x)$.
Since the integral $Z(x,\phi)$ is usually intractable in practice, APT proposes the use of ‘atomic’ proposals, APT with such proposal is known as atomic APT. Specifically, they assume that $\tilde{p}(\theta)=U_{\Theta}$, where $U_{\Theta}$ is a uniform distribution over a finite set $\Theta=\{\theta_{1},\dots,\theta_{M}\}$. The uniform setting of the proposal distribution allows for the analytical computation of $Z(x,\phi)$. With this method, (3) and (5) can be reformulated as
$\tilde{p}(\theta|x)=\frac{p(\theta|x)/p(\theta)}{\sum_{\theta^{\prime}\in\Theta}p(\theta^{\prime}|x)/p(\theta^{\prime})},\quad\tilde{q}_{x,\phi}(\theta)=\frac{q_{F(x,\phi)}(\theta)/p(\theta)}{\sum_{\theta^{\prime}\in\Theta}q_{F(x,\phi)}(\theta^{\prime})/p(\theta^{\prime})}.$ (6)
Proposition 1 in [25] provides the consistency guarantees of atomic APT: given that each $\Theta$ is constructed by sampling $\theta^{\prime}$ from a distribution that covers the target $p(\theta|x_{o})$, atomic APT is able to recover the full posterior.
However, to the best of our knowledge, the use of atomic proposals makes it challenging to analyze its convergence behaviour. Therefore unable to explain the unexpected low performance in some tasks [13] with the existing convergence results [6], both in terms of insight and theory. As an alternative approach to estimate $Z(x,\phi)$, the nested estimation, which enjoys a comprehensive theoretical framework and comparable performance, is studied in the next section.
2.2 Nested APT method
For ease of presentation, we denote $g_{\phi}(x,\theta):=q_{F(x,\phi)}(\theta)/p(\theta)$. We reformulate (4) as
@@ 2.2 Nested APT method
$=-\mathbb{E}_{\tilde{p}(\theta,x)}\left[\log q_{F(x,\phi)}(\theta)+\log\frac{\tilde{p}(\theta)}{p(\theta)}\right]+\mathbb{E}_{\tilde{p}(x)}[\log Z(x,\phi)]$
@@
:=~$:=\tilde{\mathcal{L}}(\phi)-\mathbb{E}_{\tilde{p}(\theta)}\left[\log\tilde{p}(\theta)\right],$
@@
~$\tilde{\mathcal{L}}(\phi)=-\mathbb{E}_{\tilde{p}(\theta,x)}\left[\log g_{\phi}(x,\theta)\right]+\mathbb{E}_{\tilde{p}(x)}\left[\log\mathbb{E}_{\tilde{p}(\theta^{\prime})}\left[g_{\phi}(x,\theta^{\prime})\right]\right],$ (7)
is selected as the loss function, $-\mathbb{E}_{\tilde{p}(\theta)}\left[\log\tilde{p}(\theta)\right]$ in the last column is dropped for it is independent of $\phi$. We employ stochastic gradient methods to optimize the loss function in this paper, therefore $\nabla_{\phi}\tilde{\mathcal{L}}(\phi)$ is of interest. By allowing the interchange of expectation and gradient operator, the gradient of (7) is given by
$\nabla_{\phi}\tilde{\mathcal{L}}(\phi)=-\mathbb{E}_{\tilde{p}(\theta,x)}\left[\nabla_{\phi}\log g_{\phi}(x,\theta)\right]+\mathbb{E}_{\tilde{p}(x)}\left[\nabla_{\phi}\log\mathbb{E}_{\tilde{p}(\theta^{\prime})}\left[g_{\phi}(x,\theta^{\prime})\right]\right].$ (8)
Notice that (8) holds a similar formulation of the gradient use in the optimization procedure of Bayesian experimental design [9, 16, 34, 37]. Given any query point $(\theta,x)\sim\tilde{p}(\theta,x)$, we denote the corresponding query of the loss and gradient as
$\psi_{\phi}$ $=-\log g_{\phi}(x,\theta)+\log Z(x,\phi),$ (9)
$\rho_{\phi}$ $=-\nabla_{\phi}\log g_{\phi}(x,\theta)+\nabla_{\phi}\log Z(x,\phi),$ (10)
so that $\mathbb{E}_{\tilde{p}(\theta,x)}[\psi_{\phi}]=\tilde{\mathcal{L}}(\phi)$ and $\mathbb{E}_{\tilde{p}(\theta,x)}[\rho_{\phi}]=\nabla_{\phi}\tilde{\mathcal{L}}(\phi)$. With the abuse of notation aiming at easing representation, unless otherwise specified in the following paper, $\mathcal{L}(\phi)$ and $\nabla_{\phi}\mathcal{L}(\phi)$ will be used to represent (7) and (8) respectively.
@@
Nested estimators for the loss function (7) and its gradient (8) are then given by the mean of $N$ iid copies of their queries
@@
where $\theta_{ij}^{\prime}\stackrel{{\scriptstyle iid}}{{\sim}}\tilde{p}(\theta^{\prime})$ constitute the inner sample with a size of $M$ and $(\theta_{i},x_{i})\stackrel{{\scriptstyle iid}}{{\sim}}\tilde{p}(\theta,x)$ constitute the outer sample with a size of $N$, $\psi^{(i)}_{\phi,M},\ \rho^{(i)}_{\phi,M}$ are iid copies of $\psi_{\phi,M}$ and $\rho_{\phi,M}$ respectively. Due to the nonlinearity of the logarithm, (14) is a biased estimator. Using similar arguments in [53], it is not difficult to evaluate that the order of variance of the estimator of nested APT ^$\hat{{\mathcal{L}}}^{\mathrm{Ne}}(\phi)$ is of $O(1/N)$, and $O(1/M)$ for its bias.
2.3 Performance evaluations for comparing nested APT with atomic APT
In this section, we conduct a series of experiments to demonstrate the performance of nested APT compared to atomic APT. For detailed experimental setting, we refer the readers to Section 5. The results are reported in Figure 1, from which we observe that the proposed nested APT achieves comparable performance with atomic APT. A noteworthy advantage of our proposed nested APT method is that it enables the application of a series of existing result for an optimizer with a biased gradient.
Due to its biased nature, the nested APT method may not directly benefit from the existing variance reduction techniques, as the effects of variance and bias on the optimal gap are ‘tangled’. To address this issue, we propose to make use of the existing multi-level Monte Carlo (MLMC) techniques for the improvement of the nested estimator.
A
B
C
Figure 1: Density plot for nested APT and atomic APT. A. Two-moon model, from left to right: available ground truth, atomic APT with inner samples $M=100$, nested APT with inner samples $M=100$. B. Lotka-Volterra model, from left to right: ground truth simulated with SMC-ABC [4], atomic APT with $M=100$, nested APT with $M=100$. C. M/G/1 queue model, the setting are the same with Lotka-Volterra.
3 Unbiased multilevel nested APT

### v2

@@ Abstract
There has been a growing interest in studying sequential neural posterior estimation (SNPE) techniques for their advantages in dealing with simulation-based models with intractable likelihoods. They are devoted to learning the posterior from adaptively proposed simulations using neural network-based conditional density estimators. As a SNPE technique, the automatic posterior transformation (APT) method proposed by Greenberg et al. (2019) performs notably and scales to high dimensional data. However, the APT method bears the computation of an expectation of the logarithm of an intractable normalizing constant, i.e., a nested expectation. Although atomic APT was proposed to solve this by discretizing the normalizing constant, it remains challenging to analyze the convergence of learning. In this paper, we propose a nested APT method to estimate the involved nested expectation instead. This facilitates establishing the convergence analysis. Since the nested estimators for the loss function and its gradient are biased, we make use of unbiased multi-level Monte Carlo (MLMC) estimators for debiasing. To further reduce the excessive variance of the unbiased estimators, this paper also develops some truncated MLMC estimators by taking account of the trade-off between the bias and the average cost. Numerical experiments for approximating complex posteriors with multimodal in moderate dimensions are provided.
@@ 1 Introduction
Simulator-based models are widely used across various scientific disciplines, including neuroscience [46], physics [7, 23], biology [28, 38, 45] , and inverse graphics [52]. These models serve as crucial tools for describing and comprehending investigated processes based on observed data. However, when applying traditional Bayesian inference to simulator-based models, challenges arise, such as the intractable likelihood function $p(x|\theta)$ and the computational expense associated with running the simulator.
@@
Posterior density estimation approaches approximate the posterior of interest $p(\theta|x_{o})$ with a family of density estimators $q_{\phi}(\theta)$, where $\phi$ is the parameters of the density estimators. Optimization-based approaches are widely used in these methods, and the Kullback-Leibler (KL) divergence between $p(\theta|x)$ and $q_{\phi}(\theta)$, which measures the differences between two densities, is commonly chosen as the loss function. Variational Bayes (VB), as a computationally optimization-based effective method for approximating the posterior distribution of a Bayesian problem, is widely used. In the likelihood-free context, [57] developed a new VB method with an intractable likelihood, while [29] proposed an unbiased VB method based on nested MLMC. However, these methods tend to fail in cases where simulations are expensive, as the nested estimation in these methods requires additional simulation procedures.
For this reason, there has been a growing interest in employing neural networks to represent probability density recently, particularly normalizing flows [37]. When the inference problem is focused solely on the observation $x_{o}$, the data efficiency can be improved using the sequential training schemes proposed in the sequential neural posterior estimation (SNPE) methods [24, 41, 47]. In this approach, model parameters are drawn from a proposal distribution that is more informative about $x_{o}$ compared to the prior distribution. However, SNPE requires a modification of the loss function compared to neural posterior estimation (NPE) to ensure that the neural network approximates the true posterior $p(\theta|x)$. Many approaches have been developed to address this issue.
Among the SNPE methods, automatic posterior transformation (APT) requires the computation of the expectation of the logarithm of an intractable normalizing constant, which is a nested expectation. They then propose to use the atomic APT to discretize the normalizing constant. However, it remains challenging to utilize the existing analysis techniques [6] to perform convergence analysis to the best of our knowledge. To address this limitation, we propose the nested APT method, a method enable the convergence analysis without sacrificing the performance. We conduct a thorough analysis of the bias and variance of the loss estimators, moreover, we provide a convergence result for this biased estimator in the case of stochastic gradient descent (SGD). To evaluate the effectiveness of the nested APT, we conduct a series of comprehensive numerical experiments.
@@
The remainder of this paper is organized as follows. In Section 2, we propose the nested APT method and provide insights into possible failures when directly using the nested APT. Moreover, we propose a strategy to address this failure. We also present the experimental results and analyze them in this section. In Section 3, we introduce the basics of MLMC methods, including their formulation and theoretical analysis of the random sequence utilized in MLMC methods. Equipped with these tools, we develop both unbiased MLMC methods and truncated MLMC methods and analyze the order of variance for the losses and gradients, as well as the average cost. In Section 4, we provide convergence analysis for nested APT and truncated MLMC methods in the case of SGD. In  Section 5, we conduct a series of numerical experiments on benchmark tasks. The results of these experiments are presented and analyzed. Finally, we conclude this paper with some remarks in Section 6. A lengthy proof and some numerical results are deferred in the appendices.
@@ 2.1 Problem formulation
Assume that the prior distribution has a density $p(\theta)$ with respect to the Lebesgue measure, where the model parameter of interest $\theta\in\Theta$. We focus on the finite-dimensional parameter space $\Theta$. Given an observed sample $x_{o}$, our objective is to perform inference on the posterior $p(\theta|x_{o})\propto p(\theta)p(x_{o}|\theta)$. However, in many cases, the likelihood $p(x|\theta)$ either lacks an explicit expression or presents challenges in direct evaluation. Instead, it can be expressed using a ‘simulator’ approach, wherein, given a fixed model parameter $\theta$, we can sample $x$ from $p(x|\theta)$.
@@
which is not less than 00 by using Jensen’s inequality and attains its minima when $q(\theta)$ agrees with $p(\theta)$, making it suitable as a loss function. When the likelihood $p(x|\theta)$ is tractable, one can directly approximate $p(\theta|x_{o})$ by minimizing the KL divergence between the target distribution and the proposed estimator $q(\theta)$ within a certain family of distributions. In the likelihood-free context, this can be viewed as a problem of conditional density estimation. Within this framework, a conditional density estimator $q_{F(x,\phi)}(\theta)$ based on a neural network [47, 24], is utilized to approximate $p(\theta|x)$ over the admissible set of tuning parameter $\phi\in\Phi$, where $\Phi\subset\mathbb{R}^{d}$ is the space of the neural network parameter with a finite dimension $d$. To this end, we focus on minimizing the following average KL divergence under the marginal distribution with density $p(x)=\int p(\theta)p(x|\theta)\mathrm{d}\theta$
@@
$:=L_{1}(\phi)+\iint p(\theta,x)\log p(\theta|x)\mathrm{d}x\mathrm{d}\theta,$
@@
$L_{1}(\phi):=-\mathbb{E}_{p(\theta,x)}\left[\log q_{F(x,\phi)}(\theta)\right],$ (1)
is used as the loss function. However, since $L_{1}(\phi)$ is intractable, we use its empirical estimator instead
$\hat{L}_{1}(\phi)=-\frac{1}{N}\sum_{i=1}^{N}\log q_{F(x_{i},\phi)}(\theta_{i}),$ (2)
@@
Since we aim to conduct conditional density estimation at $x_{o}$, it is essential to utilize a proposal $\tilde{p}(\theta)$ that provides more informative priors regarding $x_{o}$ in comparison to the prior $p(\theta)$. After initializing $\tilde{p}(\theta)$ as $p(\theta)$, we then want the approximation of $p(\theta|x_{o})$ to serve as a good candidate of the proposal in the following simulation. This conditional density estimation with an adaptively chosen proposal is called sequential neural posterior estimation (SNPE). However, after replacing $p(\theta,x)$ with $\tilde{p}(\theta,x)=\tilde{p}(\theta)p(x|\theta)$ in (1), it is observed that $q_{F(x,\phi)}(\theta)$ approximates the proposal posterior
$\tilde{p}(\theta|x):=p(\theta|x)\frac{\tilde{p}(\theta)p(x)}{p(\theta)\tilde{p}(x)},$ (3)
where $\tilde{p}(x)=\int\tilde{p}(\theta)p(x|\theta)\mathrm{d}\theta$. Hence, we need to adjust the loss function to make $q_{F(x,\phi)}(\theta)$ approximate the true posterior $p(\theta|x)$.
@@
$L_{2}(\phi):=\mathbb{E}_{\tilde{p}(\theta,x)}\left[-\log\tilde{q}_{F(x,\phi)}(\theta)\right],$ (4)
@@
$\tilde{q}_{F(x,\phi)}(\theta)=q_{F(x,\phi)}(\theta)\frac{\tilde{p}(\theta)}{p(\theta)}\frac{1}{Z(x,\phi)},\quad Z(x,\phi)=\int\frac{q_{F(x,\phi)}(\theta^{\prime})}{p(\theta^{\prime})}\tilde{p}(\theta^{\prime})\mathrm{d}\theta^{\prime},$ (5)
$Z(x,\phi)$ here denotes the normalizing constant. Proposition 1 in [47] shows that if $q_{F(x,\phi)}(\theta)$ is expressive enough that $\tilde{q}_{F(x,\phi^{*})}(\theta)=\tilde{q}(\theta)$ for some parameter $\phi^{*}$, then $q_{F(x,\phi^{*})}(\theta)=q(\theta|x)$. SNPE methods like APT iteratively refine the parameter $\phi$ and proposal $\tilde{p}(\theta)$ over a series of iterations, commonly called the ‘rounds’ of training. In the $r$-th round, a distinct proposal $\tilde{p}_{r}(\theta)$ is used, leading to a different loss function.
Since the integral $Z(x,\phi)$ is usually intractable in practice, APT proposes the use of ‘atomic’ proposals. APT with such proposals is known as atomic APT. Specifically, they assume that $\tilde{p}(\theta)$ is a discrete uniform distribution $\mathcal{U}\{\theta_{1},\dots,\theta_{M}\}$, where $\theta_{i}$ are sampled from some distribution. The uniform setting of the proposal distribution allows for the analytical computation of $Z(x,\phi)$. With this method, (3) and (5) can be reformulated as
$\tilde{p}(\theta|x)=\frac{p(\theta|x)/p(\theta)}{\sum_{i=1}^{M}p(\theta_{i}|x)/p(\theta_{i})},\quad\tilde{q}_{F(x,\phi)}(\theta)=\frac{q_{F(x,\phi)}(\theta)/p(\theta)}{\sum_{i=1}^{M}q_{F(x,\phi)}(\theta_{i})/p(\theta_{i})}.$ (6)
Proposition 1 in [24] provides the consistency guarantees of atomic APT: given that $\theta_{1},\dots,\theta_{M}$ are generated from a distribution that covers the target $p(\theta|x_{o})$ support, atomic APT can recover the full posterior.
However, to the best of our knowledge, the use of atomic proposals makes it challenging to analyze their convergence behavior. Therefore we are unable to explain the unexpected low performance in some tasks [12] with the existing convergence results [6], both in terms of insight and theory. As an alternative approach to estimating $Z(x,\phi)$, the nested estimation, which enjoys a comprehensive theoretical framework and comparable performance, is studied in the next section.
2.2 Nested estimation
For ease of presentation, we denote
$g_{\phi}(x,\theta):=\frac{q_{F(x,\phi)}(\theta)}{p(\theta)}.$
We reformulate (4) as
@@ 2.2 Nested estimation
$:=\mathcal{L}(\phi)-\mathbb{E}_{\tilde{p}(\theta)}\left[\log\tilde{p}(\theta)\right],$ (7)
@@
$\mathcal{L}(\phi):=-\mathbb{E}_{\tilde{p}(\theta,x)}\left[\log g_{\phi}(x,\theta)\right]+\mathbb{E}_{\tilde{p}(x)}\left[\log\mathbb{E}_{\tilde{p}(\theta^{\prime})}\left[g_{\phi}(x,\theta^{\prime})\right]\right]$ (8)
is selected as the loss function, and the last term $-\mathbb{E}_{\tilde{p}(\theta)}\left[\log\tilde{p}(\theta)\right]$ in (7) is dropped for it is independent of $\phi$. We employ stochastic gradient methods to optimize the loss function in this paper, and therefore $\nabla_{\phi}\mathcal{L}(\phi)$ is of interest. By allowing the interchange of expectation $\mathbb{E}[\cdot]$ and gradient operator $\nabla_{\phi}$, the gradient of (8) is given by
$\nabla_{\phi}\mathcal{L}(\phi):=-\mathbb{E}_{\tilde{p}(\theta,x)}\left[\nabla_{\phi}\log g_{\phi}(x,\theta)\right]+\mathbb{E}_{\tilde{p}(x)}\left[\nabla_{\phi}\log\mathbb{E}_{\tilde{p}(\theta^{\prime})}\left[g_{\phi}(x,\theta^{\prime})\right]\right].$ (9)
Notice that (9) holds a similar formulation of the gradient used in the optimization procedure of Bayesian experimental design [9, 15, 33, 36]. Given any query point $(\theta,x)\sim\tilde{p}(\theta,x)$, we denote the corresponding query of the loss and gradient as
$\psi_{\phi}$ $:=-\log g_{\phi}(x,\theta)+\log Z(x,\phi),$ (10)
$\rho_{\phi}$ $:=-\nabla_{\phi}\log g_{\phi}(x,\theta)+\nabla_{\phi}\log Z(x,\phi),$ (11)
so that $\mathbb{E}_{\tilde{p}(\theta,x)}[\psi_{\phi}]=\mathcal{L}(\phi)$ and $\mathbb{E}_{\tilde{p}(\theta,x)}[\rho_{\phi}]=\nabla_{\phi}\mathcal{L}(\phi)$.
@@
Nested estimators for the loss function (8) and its gradient (9) are then given by the mean of $N$ independent identically distributed (iid) copies of their queries
@@
where $\theta_{ij}^{\prime}\stackrel{{\scriptstyle iid}}{{\sim}}\tilde{p}(\theta^{\prime})$ constitute the inner sample with a size of $M$ and $(\theta_{i},x_{i})\stackrel{{\scriptstyle iid}}{{\sim}}\tilde{p}(\theta,x)$ constitute the outer sample with a size of $N$, $\psi^{(i)}_{\phi,M},\ \rho^{(i)}_{\phi,M}$ are iid copies of $\psi_{\phi,M}$ and $\rho_{\phi,M}$ respectively. Due to the nonlinearity of the logarithm, (15) is a biased estimator. Using similar arguments in [53], it is not difficult to evaluate that the nested estimator ^$\hat{{\mathcal{L}}}^{\mathrm{Ne}}(\phi)$ has a variance of $\mathcal{O}(1/N)$ and a bias of $\mathcal{O}(1/M)$. For a given mean squared error (MSE) $\epsilon$, choosing $N=\mathcal{O}(\epsilon^{-2})$ and $M=\mathcal{O}(\epsilon^{-1})$ yields a computational complexity of $\mathcal{O}(NM)=\mathcal{O}(\epsilon^{-3})$.
Compared to nested APT, the bias properties of atomic APT are less straightforward to analyze theoretically. However, empirical results show that the performance of both methods is similar. Visual comparisons of their approximated posteriors are provided in Appendix C. However, since nested APT updates the model with a biased gradient, the update direction will systematically deviate from the optimal solution of the loss function, potentially preventing the model from converging to the global optimum or leading it to an incorrect local extremum, increasing training time and computational cost. What’s worse, the parameter updates may fail to reflect the characteristics of the target posterior. To address this issue, we propose to make use of the existing MLMC techniques for the improvement of the nested estimator.
3 Multilevel nested APT